OpenStudy (anonymous):

If Dr. Suzuki rides his bike to his office, he averages 6 miles per hour. If he drives his car, he averages 18 miles per hour. His time driving is 1/2 hour less than his time bicycling. How far is his office?

OpenStudy (dan815):

dr suzuki should drive faster

OpenStudy (dan815):

v=d/t

OpenStudy (dan815):

6*t+0.5=d 18*t=d

OpenStudy (dan815):

solve for d, 2 equations 2 unknowns, linearly independant

OpenStudy (dan815):

oh sorry umm its 6(t+0.5)=d

OpenStudy (whpalmer4):

Alternative formulation: let \(x\) be distance to office: Time on bike = \(x/6\) Time in car = \(x/18\) Time in car = time on bike - 1/2 \[\frac{x}{18} = \frac{x}{6}-\frac{1}{2}\]Multiply by 18 to clear fractions \[x = 3x - 9\]

OpenStudy (popeyes):

Speed of bike (b) = 6 miles/hr Speed of car (c) = 18 miles/hr Distance between home to office = d Formula: Time = Distance/Speed Time taken by bike = d/6 Time taken by car = d/18 Car takes 1/2 hours less than a bike. \[\frac{ d }{ 6 }- \frac{ d }{ 18 }=\frac{ 1 }{ 2 } \\ \\ \frac{ 3d }{ 18 }- \frac{ d }{ 18 }= \frac{ 1 }{ 2 }\\ \\ \frac{ 2 }{ d }= \frac{ 1 }{ 2 } \\ \\ 2d = 9 \\ \\ d = 9/2 \ miles\] So, Dr. Suzuki's office is 9/2 (or 4.5) miles away.